(a) Show that sin(3x) + sin(x) = 4sin(x)cos^2(x)

(b) Find all the angles between 0 and "\pi" which satisfy the equation sin(3x) + sin(x) = 2(cos^2)(x).


1
Expert's answer
2021-10-18T07:41:21-0400

(a) Let us show that "\sin(3x) + \sin(x) = 4\sin(x)\cos^2(x)." Indeed,


"\sin(3x) + \sin(x)=2\sin (\frac{3x+x}2)\cos (\frac{3x-x}2)=2\sin (2x)\cos (x)\\=2\cdot2\sin(x)\cos(x)\cos(x)=4\sin(x)\cos^2(x)."


(b) Let us find all the angles between 0 and "\pi" which satisfy the equation "\sin(3x) + \sin(x) = 2\cos^2(x)," which is equivalent to "4\sin(x)\cos^2(x)=2\cos^2(x)," and hence to "2\cos^2(x)(2\sin(x)-1)=0." It follows that "\cos(x)=0" or "\sin(x)=\frac{1}2," and hence "x=\frac{\pi}2+\pi n" or "x=\frac{\pi}6+2\pi n" or "x=\frac{5\pi}6+2\pi n, \ n\in\Z." Therefore, the angles between 0 and "\pi" which satisfy the equation are "\frac{\pi}2,\ \frac{\pi}6,\ \frac{5\pi}6."